OpenStudy (anonymous):

Full electron configuration for.. 1. Copper 2. Iodine 3. Potassium 4. Bismuth 5. Zicronium

OpenStudy (anonymous):

Hey there! So how I like to think about answering this question is to first look at the definition and then think about is the number of electrons in each atom. Yay definitions! So... Electron configuration is defined as the arrangement of electrons of an atom, a molecule, or other physical structure. It concerns the way electrons can be distributed in the orbitals of the given system (atomic or molecular for instance). Do you remember Aufbau's Principle (see below in the attachments). Aufbau's principle is key in understanding electron configuration so MEMORIZE it. It can be your best friend. So let's look at the problem. 1. Copper has 29 electrons. {Note the number's after each letter's are written as exponents) a) Short hand way using a Noble gas: [Ar] 4s1 3d10 b)Full hand way: 1s2 2s2 2p6 3s2 3p6 4s1 3d10 To check what you do is add the exponents up: 2+2+6+2+6+1+10=29 2. Iodine has 53 electrons- Oh dear this one is going to be a long one.. a) Short hand way using noble gas: [Kr] 5s2 4d10 5p5 b)Full hand way: 1s2 2s2 2p6 3s2 3p6 4s2 3d10 4p6 5s2 4d10 5p5 Checking time: 2+2+6+2+6+2+10+6+2+10+5= 53 3.Potassium has 19 electrons a) Short hand: [Ar] 4s1 b) Full: 1s2 2s2 2p6 3s2 3p6 4s1 Check time: 2+2+6+2+6+1 =19 4. Bismuth has 83 electrons.. [I don't like this one already but here we go..] a)short hand: [Xe] 4f14 6s2 5d10 6p3 b) long hand: 1s2 2s2 2p6 3s2 3p6 4s2 3d10 4p6 5s2 4d10 4f14 5p6 6s2 5d10 6p3 Check: 83 5. Zicronium has 40 electrons a) short hand: [Kr] 5s2 4d2 b) Long hand: 1s2 2s2 2p6 3s2 3p6 4s2 3d10 4p6 5s2 4d2 Check: 2+2+6+2+6+2+10+6+2+2=40 SO to make your life easier, MEMORIZE Aufbau's principle. Hope that this helps! Let me know if you need more explanations.